OpenStudy (anonymous):

Which of these is a simplified form of the equation 6y + 4 = 8 + 2y + 2y? 6y = 8 10y = 12 2y = 4 4y = 12

OpenStudy (anonymous):

Combine all the like terms on both sides of the equation.

OpenStudy (anonymous):

Can you combine it to this?\[6y + 4 = 8 +4y\]

OpenStudy (anonymous):

thanks what would i get

OpenStudy (anonymous):

no u cant

OpenStudy (anonymous):

Yea, now can you put the y's on one side and the constants on the other?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think u can

OpenStudy (anonymous):

Sure, and what would you get?

OpenStudy (anonymous):

im not sure ;(

OpenStudy (anonymous):

When you move a term over the equality sign, you just negate the value.

OpenStudy (anonymous):

For example, \[x + 2y = 3\] is equivalent to \[x = 3 - 2y\]

OpenStudy (anonymous):

Can you try moving the terms over to the other side?

OpenStudy (anonymous):

Notice how in my example, 2y on the left hand side becomes -2y on the right hand side.

OpenStudy (anonymous):

so C? :)

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

THANKSSS can u help with another?